(3 points) The following table lists the proportion of employed people by each major industry in Japan as of December 2012. This includes all of the industries Major Industry Proportion Agriculture and forestry Construction Manufacturing Information and communications Transport and postal activities Wholesale and retail trade Scientific, professional, technical Accommodations Personal services Education Medical Services Government 0.0328 0.0842 0.1722 0.0330 0.0585 0.1786 0.0362 0.0666 0.0412 0.0511 0.1273 0.0797 0.0386 Suppose a major industry is randomly selected. Consider the events A-Construction or Manufacturing) B not Government C(Medical Find the following probabilities (0.3 pts.) a) P(A) Submit Answer Tries 0/5 (0.3 pts.) b) P(B) Submit Answ Tries 0/5 (0.5 pts.) c) P(C) Submit Answer Tries 0/5 (0.3 pts.) d) P(A n B) Submit Answer Tries 0/5 (0.3 pts.) e) P(A U C) Submit Answ Tries 0/5 (0.5 pts.) f) P(AnC) Submit Answer Tries 0/5 (0.8 pts: 0.2 for parts a), b), d), e)) Please show your work for the above questions here
Answer #1

a)P(A)=0.0842+0.1722=0.2564

b)P(B)=1-0.0386=0.9614

c)P(C)=0.1273

d)P(A n B)=0.2564

e)P(A u C)=0.2564+0.1273=0.3837

f)P(A n C)=0